OpenStudy (anonymous):

Which of the following is NOT equivalent to ? A. pie 4/9 B. 0.6 continuos C. 0.23 D. 4/6

OpenStudy (anonymous):

equivalent to what?

OpenStudy (anonymous):

oh sorry :b ... 2/3

OpenStudy (hoblos):

i think you mean sqrt4/9 in A sqrt4/9 = 4/6 = 0.666 = 2/3 =>the answer is C

OpenStudy (anonymous):

Thanks (:

OpenStudy (hoblos):

my pleasure XD